I am thinking of three consecutive numbers. If I multiply the first with the second and then subtract three times the third, the result is 57. What are the numbers? Can someone please show me the steps to get to the answer? Thank you.
Answer by edjones(8007)

Let the numbers be x, x+1 and x+2.
x(x+1)-3(x+2)=57
x^2+x-3x-6=57
x^2-2x-6=57
x^2-2x-63=0
(x-9)(x+7)=0
x=9, x=-7 Two different sets of correct answers: 9, 10, 11 and -7, -6, -5
.
Ed